# Multi-Modal Hypergraph Enhanced LLM Learning for Recommendation

### v1

@@ Abstract.
The burgeoning presence of Large Language Models (LLM) is propelling the development of personalized recommender systems. Most existing LLM-based methods fail to sufficiently explore the multi-view graph structure correlations inherent in recommendation scenarios. To this end, we propose a novel framework, Hypergraph Enhanced LLM Learning for multimodal Recommendation (HeLLM), designed to equip LLMs with the capability to capture intricate higher-order semantic correlations by fusing graph-level contextual signals with sequence-level behavioral patterns. In the recommender pre-training phase, we design a user hypergraph to uncover shared interest preferences among users and an item hypergraph to capture correlations within multimodal similarities among items. The hypergraph convolution and synergistic contrastive learning mechanism are introduced to enhance the distinguishability of learned representations. In the LLM fine-tuning phase, we inject the learned graph-structured embeddings directly into the LLM’s architecture and integrate sequential features capturing each user’s chronological behavior. This process enables hypergraphs to leverage graph-structured information as global context, enhancing the LLM’s ability to perceive complex relational patterns and integrate multimodal information, while also modeling local temporal dynamics. Extensive experiments demonstrate the superiority of our proposed method over state-of-the-art baselines, confirming the advantages of fusing hypergraph-based context with sequential user behavior in LLMs for recommendation.
1. Introduction
@@
Recommender systems aim to mitigate information overload by uncovering and satisfying users’ implicit preferences and needs (Ying et al., 2018; He et al., 2020). Large Language Models (LLM) have demonstrated unprecedented capabilities across various domains like NLP (Achiam et al., 2023), Computer Vision (Wang et al., 2024; Li et al., 2023b), and Medicine (Thirunavukarasu et al., 2023), and recent efforts have explored their potential to enhance the personalization of recommendation.
@@
Recent studies (Guo et al., 2023; Wu et al., 2024a; Li et al., 2023c; Zhang et al., 2023b; Liao et al., 2024; Lin et al., 2024; Wu et al., 2024b) employ Parameter-Efficient Fine-Tuning (PEFT) (Ding et al., 2023; Lialin et al., 2023) to bridge the knowledge gap between LLM and recommender models by optimizing selective or additional lightweight parameters. These models are typically structured within a Recommendation Embedding Schema, consisting of a pre-trained lightweight recommendation backbone that learns embeddings for users and items, a linear mapping layer that aligns these embeddings with the LLM’s representation space, and an LLM backend with textual prompt for enriching semantic understanding.
@@
As shown in Fig. 1, existing methods fall into two main categories: the first (Zhang et al., 2023b) represents whole user-item interactions as a bipartite graph, utilizing Graph Neural Network (GNN) backbones to capture Collaborative Filtering (CF) signals; the second (Li et al., 2023a; Liao et al., 2024) models each user’s interaction history with a time-series backbone to capture user’s sequential behavior patterns. For example, Collm (Zhang et al., 2023b) integrates learned collaborative representations of LightGCN (He et al., 2020) into LLM for recommendation. LLaRA (Liao et al., 2024) harness traditional sequential recommenders such as GRU4Rec (Hidasi, 2015) and SASRec (Kang and McAuley, 2018) to encode “sequential behaviors of users” as a distinct modality beyond texts. These methods have shown promising results in integrating user interaction patterns with LLM-driven semantics.
@@
However, in practical scenarios, both GNN-based and time-series approaches are essential, each capturing unique facets of user behavior that are indispensable for the comprehensive recommendation. GNN-based CF methods (Wang et al., 2019b; He et al., 2020; Wu et al., 2022) explicitly model intricate relational structures across the entire user-item interaction space, capturing multi-hop dependencies among users and items via message propagation, and providing a global perspective on interactions. But they are limited in capturing the fine-grained temporal evolution, which is crucial for modeling users’ time-sensitive sequential behaviors. Conversely, sequence-based methods (Kang and McAuley, 2018; Zhou et al., 2020; Wu et al., 2024b) capture temporal dependencies yet fail to adequately mine abundant structured correlations (e.g., social networks) among users and items within the recommendation context, lacking a global CF supervisory signal.
@@
Evidently, LLM-based methods hold the potential to improve performance by increasing the power of pre-trained representations. Despite the complementary advantages of CF-based and sequence-based embeddings, their mutual effect remains largely unexplored. Existing LLM-based methods (Liao et al., 2024; Zhang et al., 2023b) primarily focus on learnable ID attributes and user-item bipartite graphs, also with limited exploration of high-order shared user preferences and multimodal side information correlations among items. Typically, these methods leverage enhanced prompt learning (Lester et al., 2021) (tied to pre-trained embeddings) and additional LoRA (Hu et al., 2021) parameters (independent of pre-trained embeddings) for PEFT of LLM. Recent study (Ding et al., 2023) indicates that fine-tuning intermediate layers (e.g., via LoRA) yields superior performance compared to adjustments in the initial (prompt learning-like) or final layers. This insight motivates us to inject rich graph-structured information into LLM internals during fine-tuning, aiming to unlock graph-perceived potential. In summary, we identify two core objectives for advancing LLM-based recommender systems: (1) enhancing the expressive power of pre-trained representations, and (2) effectively integrating rich graph structure information to strengthen the LLM’s ability to perceive relational patterns.
@@
2. Related Work
2.1. GNN-based Recommender Systems.
@@ 2. Related Work
Graph neural networks (GNN) are widely used in recommender systems to model graph-structured data and learn representations (Wu et al., 2022). These methods generally fall into five categories: 1) Collaborative filtering based on user-item interactions (e.g., NGCF (Wang et al., 2019b), LightGCN (He et al., 2020)). 2) Sequential and session recommendation based on time-series user historical behaviors (e.g., SR-GNN (Wu et al., 2019), GC-SAN (Xu et al., 2019)). 3) Social recommendation based on user associations (e.g., ESRF (Yu et al., 2020), GraphRec (Fan et al., 2019)). 4) Knowledge graph based recommendation (Wang et al., 2019c, a). 5) Multimodal based recommendation (He and McAuley, 2016; Wei et al., 2019, 2020; Zhang et al., 2021; Wang et al., 2021; Yi and Chen, 2021; Zhou and Shen, 2023; Chen et al., 2017, 2019; Wei et al., 2023; Zhou et al., 2023; Tao et al., 2022). In contrast to prior approaches, our method explores second-order semantic correlations at the user and item levels to enhance multimodal recommendation and enrich LLM capabilities to perceive and reason over complex graph structures.
2.2. Parameter-Efficient Fine-Tuning.
@@
Parameter-Efficient Fine-Tuning methods (Ding et al., 2023; Lialin et al., 2023) only optimize a small subset of parameters or introduce additional ones in the LLM while keeping the majority frozen to cut down computational and storage costs compared with full-tuning. These methods fall into four categories: 1) Addtion-based methods, which introduce extra learnable parameters into LLM and only train them (e.g. Adapter (Houlsby et al., 2019), Prompt-Tuning (Lester et al., 2021), Prefix-Tuning (Li and Liang, 2021), Ladder-Side Tuning (Sung et al., 2022)). 2) Selection-based methods, which selectively fine-tune a limited subset of parameters in LLM (e.g. BitFit (Zaken et al., 2021), DiffPruning (Guo et al., 2020), FishMask (Sung et al., 2021)). 3) Reparametrization-based methods, which reparameterize parameters of LLM into a low-dimensional intrinsic subspace using low-rank transformation (e.g. LoRa (Hu et al., 2021), Intrinsic SAID (Aghajanyan et al., 2020)). 4) Hybrid-based methods, which incorporate various approaches (e.g. MAM Adapter (He et al., 2021), UniPELT (Mao et al., 2021)).
2.3. Large Language Models in Recommendation.
@@
Large Language Models (Vaswani et al., 2017; Fan et al., 2023) use transformer-based deep learning architecture, combined with extensive corpus data and parameters for training, showcasing unparalleled capabilities in understanding and generating human-like language (Achiam et al., 2023; Zhao et al., 2023). Recent LLM-based recommendation methods fall into two categories: 1) Non-tuning-based methods (In-context learning methods (Hou et al., 2024)) (Gao et al., 2023; Liu et al., 2023; Zhang et al., 2023a; Dai et al., 2023; Hou et al., 2024; Wang and Lim, 2023; Ren et al., 2023; He et al., 2023; Sanner et al., 2023) which do not fine-tune the LLM and leverage the rich knowledge of the LLM for zero-shot and few-shot recommendations. 2) PEFT-based methods (Bao et al., 2023; Li et al., 2023d; Cui et al., 2022; Wu et al., 2024b; Wang et al., 2022; Guo et al., 2023; Wu et al., 2024a; Li et al., 2023c; Zhang et al., 2024) that fine-tune a small set of LMM weights or add a few extra trainable parameters to synchronize the knowledge of LLM and recommendation models. For example, TALLRec (Bao et al., 2023) fine-tunes the additional injected Low-Rank Adaptation (LoRa) (Hu et al., 2021), which employs a simple low-rank matrix decomposition to parameterize the weights in LLM. POD (Li et al., 2023d) applies prompt tuning by introducing a trainable prompt token to bridge between IDs and words. LLaRA (Liao et al., 2024) integrates sequential recommendation representation with LLMs by treating user behaviors as a distinct modality, aligning ID-based embeddings with textual features through hybrid prompting to enhance behavioral reasoning in LLMs. CoLLM (Zhang et al., 2023b) enhances LLM-based recommendation by integrating collaborative information from traditional models into LLMs’ input space via collaborative embeddings. Distinct from prior work, we explore multimodal-based multi-view graph structures to enhance the ability of representation, injecting them into LLM internals during fine-tuning to strengthen the LLM’s capacity for perceiving complex relational patterns.
3. The Proposed Method
3.1. Overview
@@
3.2. User-to-User (U2U) Hypergraph Learning
At the user level, we construct a U2U hypergraph, representing users as nodes and their item interactions as hyperedges to capture shared user preferences. This enables the exploration of second-order semantic relationships among users through hypergraph convolution, with items acting as bridges. The hypergraph encoder is constructed using stacked hypergraph convolutional layers, as in HGNN (Feng et al., 2019), to learn user-level embeddings. Formally,
@@ 3.2. User-to-User (U2U) Hypergraph Learning
where denotes the reorganized U2U hypergraph after convolution, denotes user embeddings initialized with the Xavier (Glorot and Bengio, 2010). Here, applies layer normalization (e.g., laplacian normalization) to prevent gradient explosion. The hypergraph structure facilitates a two-step information transfer: initially from nodes to hyperedges, and subsequently from hyperedges back to nodes. Specifically, in Eq. (2), gathers user information into item hyperedges, while the matrix product with integrates this hyperedge information back into user nodes. The operation quantifies the shared item connections between users. The final user-level embedding for a user , denoted as , is derived after iterations of hypergraph convolution.
3.3. Item-to-Item (I2I) Hypergraph Learning
@@
3.4. U-I Interaction Learning
The integration of user-item representations derived from a bipartite interaction graph and those from individual user and item levels serves as the output for the recommender pre-training phase. We further design a collaborative contrastive learning mechanism to enhance the discriminative power of these representations. For efficiency, LightGCN (He et al., 2020) is employed to perform convolutions on the user-item bipartite graph. Let and denote the learned ID representations of user and item , respectively. The normalized user-level and item-level representations are fused with and to enrich their information:
@@ 3.4. U-I Interaction Learning
3.5. Graph-Enhanced LLM Learning
@@
Here, represents the collection of all items’ graph features derived from the current interactions of the user . To extract graph-structured collaborative filtering signals and intricate higher-order semantic correlations from learned representations, a bilinear mapping second-order pooling similarly used in (Wang and Ji, 2020) is employed:
@@
Finally, each historical user record is converted into natural language, with item IDs replaced by concatenating with pre-trained graph-based embeddings and sequential-based embeddings from SASRec (Kang and McAuley, 2018), the same as LLaRA (Liao et al., 2024). These embeddings are then aligned with the LLM’s feature space via a multilayer perceptron. In this work, we concentrate on processing graph-based embeddings and thus omit further details on time-series embeddings.
Baseline | Sports | Beauty | Toys | |||||||||
Conventional Recommendation Methods | ||||||||||||
LightGCN (He et al., 2020) | 0.0232 | 0.0120 | 0.0411 | 0.0165 | 0.0426 | 0.0208 | 0.0669 | 0.0269 | 0.0273 | 0.0138 | 0.0409 | 0.0172 |
VBPR (He and McAuley, 2016) | 0.0298 | 0.0150 | 0.0485 | 0.0197 | 0.0436 | 0.0216 | 0.0724 | 0.0289 | 0.0376 | 0.0196 | 0.0565 | 0.0244 |
GRCN (Wei et al., 2020) | 0.0303 | 0.0152 | 0.0502 | 0.0202 | 0.0454 | 0.0225 | 0.0758 | 0.0305 | 0.0411 | 0.0214 | 0.0610 | 0.0264 |
LATTICE (Zhang et al., 2021) | 0.0310 | 0.0155 | 0.0506 | 0.0204 | 0.0468 | 0.0234 | 0.0760 | 0.0307 | 0.0459 | 0.0228 | 0.0683 | 0.0284 |
SLMRec (Tao et al., 2022) | 0.0312 | 0.0156 | 0.0507 | 0.0205 | 0.0458 | 0.0227 | 0.0754 | 0.0301 | 0.0436 | 0.0223 | 0.0675 | 0.0284 |
BM3 (Zhou et al., 2023) | 0.0316 | 0.0160 | 0.0517 | 0.0210 | 0.0446 | 0.0217 | 0.0722 | 0.0286 | 0.0402 | 0.0204 | 0.0618 | 0.0258 |
FREEDOM (Zhou and Shen, 2023) | 0.0314 | 0.0157 | 0.0507 | 0.0205 | 0.0516 | 0.0248 | 0.0805 | 0.0321 | 0.0467 | 0.0232 | 0.0700 | 0.0291 |
Large Language Models-based Methods | ||||||||||||
Collm (Zhang et al., 2023b) | 0.0349 | 0.0182 | 0.0525 | 0.0226 | 0.0681 | 0.0393 | 0.0960 | 0.0463 | 0.0652 | 0.0376 | 0.0905 | 0.0439 |
LLaRA (Liao et al., 2024) | 0.0367 | 0.0203 | 0.0562 | 0.0252 | 0.0667 | 0.0380 | 0.0925 | 0.0445 | 0.0683 | 0.0371 | 0.0944 | 0.0437 |
HeLLM (Ours) | 0.0412 | 0.0227 | 0.0617 | 0.0279 | 0.0696 | 0.0384 | 0.1021 | 0.0465 | 0.0680 | 0.0372 | 0.1027 | 0.0460 |
3.6. Objective Optimization

### v2

# HeLLM: Multi-Modal Hypergraph Enhanced LLM Learning for Recommendation

@@ Abstract
The burgeoning presence of Large Language Models (LLM) is propelling the development of personalized recommender systems. Most existing LLM-based methods fail to sufficiently explore the multi-view graph structure correlations inherent in recommendation scenarios. To this end, we propose a novel framework, Hypergraph Enhanced LLM Learning for multimodal Recommendation (HeLLM), designed to equip LLMs with the capability to capture intricate higher-order semantic correlations by fusing graph-level contextual signals with sequence-level behavioral patterns. In the recommender pre-training phase, we design a user hypergraph to uncover shared interest preferences among users and an item hypergraph to capture correlations within multimodal similarities among items. The hypergraph convolution and synergistic contrastive learning mechanism are introduced to enhance the distinguishability of learned representations. During the LLM fine-tuning phase, we directly inject the learned graph-structured embeddings into the LLM’s architecture and integrate sequential features that capture each user’s chronological behavior. This process enables hypergraphs to leverage graph-structured information as global context, enhancing the LLM’s ability to perceive complex relational patterns and integrate multimodal information, while also modeling local temporal dynamics. Extensive experiments demonstrate the superiority of our proposed method over state-of-the-art baselines, confirming the advantages of fusing hypergraph-based context with sequential user behavior in LLMs for recommendation. The code is publicly available at https://github.com/Xu107/HeLLM.
Index Terms:
Large Language Model, Hypergraph Neural Network, Multi-Modal Recommendation.I Introduction
@@
Recommender systems aim to mitigate information overload by uncovering and satisfying users’ implicit preferences and needs [1, 2]. Large Language Models (LLM) have demonstrated unprecedented capabilities across various domains like NLP [3], Computer Vision [4, 5], and Medicine [6], and recent efforts have explored their potential to enhance the personalization of recommendation.
@@
Recent studies [7, 8, 9, 10, 11, 12, 13] employ Parameter-Efficient Fine-Tuning (PEFT) [14, 15] to bridge the knowledge gap between LLM and recommender models by optimizing selective or additional lightweight parameters. These models are typically structured within a Recommendation Embedding Schema, consisting of a pre-trained lightweight recommendation backbone that learns embeddings for users and items, a linear mapping layer that aligns these embeddings with the LLM’s representation space, and an LLM backend with textual prompt for enriching semantic understanding.
@@
As shown in Fig. 1, existing methods fall into two main categories: the first [10] represents whole user-item interactions as a bipartite graph, utilizing Graph Neural Network (GNN) backbones to capture Collaborative Filtering (CF) signals; the second [16, 11] models each user’s interaction history with a time-series backbone to capture user’s sequential behavior patterns. For example, Collm [10] integrates learned collaborative representations of LightGCN [2] into LLM for recommendation. LLaRA [11] harness traditional sequential recommenders such as GRU4Rec [17] and SASRec [18] to encode “sequential behaviors of users” as a distinct modality beyond texts. These methods have shown promising results in integrating user interaction patterns with LLM-driven semantics understanding.
@@
However, in practical scenarios, both GNN-based and time-series approaches are essential, each capturing unique facets of user behavior that are indispensable for the comprehensive recommendation. GNN-based CF methods [19, 2, 20] explicitly model intricate relational structures across the entire user-item interaction space, capturing multi-hop dependencies among users and items via message propagation, and providing a global perspective on interactions. But they are limited in capturing the fine-grained temporal evolution, which is crucial for modeling users’ time-sensitive sequential behaviors. Conversely, sequence-based methods [18, 21, 13] capture temporal dependencies yet fail to adequately mine abundant structured correlations (e.g., social networks) among users and items within the recommendation context, lacking a global CF supervisory signal.
@@
Evidently, LLM-based methods hold the potential to improve performance by increasing the power of pre-trained representations. Despite the complementary advantages of CF-based and sequence-based embeddings, their mutual effect remains largely unexplored. Existing LLM-based methods [11, 10] primarily focus on learnable ID attributes and user-item bipartite graphs, also with limited exploration of high-order shared user preferences and multimodal side information correlations among items. Typically, these methods leverage enhanced prompt learning [22] (tied to pre-trained embeddings) and additional LoRA [23] parameters (independent of pre-trained embeddings) for PEFT of LLM. Recent study [14] indicates that fine-tuning intermediate layers (e.g., via LoRA) yields superior performance compared to adjustments in the initial (prompt learning-like) or final layers. This insight motivates us to inject rich graph-structured information into LLM internals during fine-tuning, aiming to unlock graph-perceived potential. In summary, we identify two core objectives for advancing LLM-based recommender systems: (1) enhancing the expressive power of pre-trained representations, and (2) effectively integrating rich graph structure information to strengthen the LLM’s ability to perceive relational patterns.
@@
II Related Work
II-A GNN-based Recommender Systems.
@@ II Related Work
Graph neural networks (GNN) are widely used in recommender systems to model graph-structured data and learn representations [20]. These methods generally fall into five categories: 1) Collaborative filtering based on user-item interactions (e.g., NGCF [19], LightGCN [2]). 2) Sequential and session recommendation based on time-series user historical behaviors (e.g., SR-GNN [24], GC-SAN [25]). 3) Social recommendation based on user associations (e.g., ESRF [26], GraphRec [27]). 4) Knowledge graph based recommendation [28, 29]. 5) Multimodal based recommendation [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. For example, VBPR [30] enhances matrix factorization by incorporating visual features alongside item ID representations. MMGCN [31] initiated the use of graph-based modeling in multimodal recommendation by applying graph convolution networks to each modality. GRCN [32] mitigates the impact of false-positive signals by detecting and pruning noisy connections within the interaction graph. In LATTICE [33], deep networks are utilized to extract implicit item structure. FREEDOM [36] fixes these learned item structures and denoises the user-item interaction graph. AlignRec [44] mitigates semantic misalignment via three-stage alignment among content, ID, and user-item features. PGL [45] constructs principal subgraphs via global- and local-aware operators to jointly exploit co-occurrence patterns and personalized signals. In contrast to prior approaches, our method explores second-order semantic correlations at the user and item levels to enhance multimodal recommendation and enrich LLM capabilities to perceive and reason over complex graph structures.
II-B Parameter-Efficient Fine-Tuning.
@@
Parameter-Efficient Fine-Tuning methods [14, 15] only optimize a small subset of parameters or introduce additional ones in the LLM while keeping the majority frozen to cut down computational and storage costs compared with full-tuning. These methods fall into four categories: 1) Addtion-based methods, which introduce extra learnable parameters into LLM and only train them (e.g. Adapter [46], Prompt-Tuning [22], Prefix-Tuning [47], Ladder-Side Tuning [48]). 2) Selection-based methods, which selectively fine-tune a limited subset of parameters in LLM (e.g. BitFit [49], DiffPruning [50], FishMask [51]). 3) Reparametrization-based methods, which reparameterize parameters of LLM into a low-dimensional intrinsic subspace using low-rank transformation (e.g. LoRa [23], Intrinsic SAID [52]). 4) Hybrid-based methods, which incorporate various approaches (e.g. MAM Adapter [53], UniPELT [54]).
II-C Large Language Models in Recommendation.
@@
Large Language Models [55, 56] use transformer-based deep learning architecture, combined with extensive corpus data and parameters for training, showcasing unparalleled capabilities in understanding and generating human-like language [3, 57]. Recent LLM-based recommendation methods fall into two categories: 1) Non-tuning-based methods (In-context learning methods [58]) [59, 60, 61, 62, 58, 63, 64, 65, 66] which do not fine-tune the LLM and leverage the rich knowledge of the LLM for zero-shot and few-shot recommendations. 2) PEFT-based methods [67, 68, 69, 13, 70, 7, 8, 9, 71] that fine-tune a small set of LMM weights or add a few extra trainable parameters to synchronize the knowledge of LLM and recommendation models. For example, TALLRec [67] fine-tunes the additional injected Low-Rank Adaptation (LoRa) [23], which employs a simple low-rank matrix decomposition to parameterize the weights in LLM. POD [68] applies prompt tuning by introducing a trainable prompt token to bridge between IDs and words. LLaRA [11] integrates sequential recommendation representation with LLMs by treating user behaviors as a distinct modality, aligning ID-based embeddings with textual features through hybrid prompting to enhance behavioral reasoning in LLMs. CoLLM [10] enhances LLM-based recommendation by integrating collaborative information from traditional models into LLMs’ input space via collaborative embeddings. Distinct from prior work, we explore multimodal-based multi-view graph structures to enhance the ability of representation, injecting them into LLM internals during fine-tuning to strengthen the LLM’s capacity for perceiving complex relational patterns.
III The Proposed Method
III-A Overview
@@
III-B User-to-User (U2U) Hypergraph Learning
At the user level, we construct a U2U hypergraph, representing users as nodes and their item interactions as hyperedges to capture shared user preferences. This enables the exploration of second-order semantic relationships among users through hypergraph convolution, with items acting as bridges. The hypergraph encoder is constructed using stacked hypergraph convolutional layers, as in HGNN [72], to learn user-level embeddings. Formally,
@@ III-B User-to-User (U2U) Hypergraph Learning
where denotes the reorganized U2U hypergraph after convolution, denotes user embeddings initialized with the Xavier [73]. Here, applies layer normalization (e.g., laplacian normalization) to prevent gradient explosion. The hypergraph structure facilitates a two-step information transfer: initially from nodes to hyperedges, and subsequently from hyperedges back to nodes. Specifically, in Eq. (2), gathers user information into item hyperedges, while the matrix product with integrates this hyperedge information back into user nodes. The operation quantifies the shared item connections between users. The final user-level embedding for a user , denoted as , is derived after iterations of hypergraph convolution.
III-C Item-to-Item (I2I) Hypergraph Learning
@@
III-D U-I Interaction Learning
The integration of user-item representations derived from a bipartite interaction graph and those from individual user and item levels serves as the output for the recommender pre-training phase. We further design a collaborative contrastive learning mechanism to enhance the discriminative power of these representations. For efficiency, LightGCN [2] is employed to perform convolutions on the user-item bipartite graph. Let and denote the learned ID representations of user and item , respectively. The normalized user-level and item-level representations are fused with and to enrich their information:
@@ III-D U-I Interaction Learning
III-E Graph-Enhanced LLM Learning
@@
Here, represents the collection of all items’ graph features derived from the current interactions of the user . To extract graph-structured collaborative filtering signals and intricate higher-order semantic correlations from learned representations, a bilinear mapping second-order pooling similarly used in [74] is employed:
@@
Finally, each historical user record is converted into natural language, with item IDs replaced by concatenating with pre-trained graph-based embeddings and sequential-based embeddings from SASRec [18], the same as LLaRA [11]. These embeddings are then aligned with the LLM’s feature space via a multilayer perceptron. In this work, we concentrate on processing graph-based embeddings and thus omit further details on time-series embeddings.
III-F Objective Optimization